OpenStudy (anonymous):

can you check this for me? The volume of a cone is three hundred cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

Radius, r = 9 km. Volume, V = 300 cu. km. Height = h V = 1/3 π r² h h = V / (1/3 π r²) Subbing known values, h = 300 / [1/3 π (9)²] h = 300 / [1/3 π (81) h = 300 / 84.823 h = 3.53

OpenStudy (anonymous):

is this correct?

OpenStudy (dan815):

i dont understand how in the ...... heck did the radius suddenly turn into 9km from 5km

OpenStudy (jhannybean):

\[\large V = \frac13 \pi r^2 h\]\[\large \frac{3V}{\pi r^2}=h\]

OpenStudy (jhannybean):

Use that formula for h...

OpenStudy (jhannybean):

\[\large V = 300 \ km^3\]\[\large r=5 \ km\]\[\large \frac{3(300)}{\pi (5)^2}=h\]

OpenStudy (jhannybean):

\[\large h = \frac{900}{25\pi}\] Now solve.

OpenStudy (jhannybean):

What do you get?..

OpenStudy (dan815):